OpenStudy (anonymous):

4x + 2y = 12 x – y = 3

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

What are you trying to find?

OpenStudy (anonymous):

is it simplify??

OpenStudy (anonymous):

x= 2 and y=-1

OpenStudy (anonymous):

Someone help me with mine :((

OpenStudy (anonymous):

Multiply the second equation by 2: 2x - 2y = 6

OpenStudy (anonymous):

add to the first equation: (4x + 2y) + (2x - 2y) = 6 + 6

OpenStudy (mathstudent55):

Write the equations one above the other. Then divide the first equation by 2. Then add the equations. The y's will get eliminated. Solve for x. Then substitute x into one of the original equations and solve for y.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Ineedanswersbadfail Was the answers i gave you right?